An art class is making a mural for their school which has a triangle drawn in the middle. The length of the bottom of the triangle is x. Another side is 3 more than four times the length of the bottom of the triangle. The last side is 5 more than the bottom of the triangle. Write and simplify an expression for the perimeter of the triangle.
What is the expression for the perimeter of the​ triangle?

Answer by rothauserc(4718)

x +(4x+3) +(x+5) = 6x +8 = 2 * (3x+4)
:
Perimeter of triangle = 2 * (3x+4)
: